SolveIt is under development
SolveItNCERT · CBSE · NEET

NCERT Exemplar · Class 12 Mathematics Probability

108 questions · 108 still being checked

EXERCISE 13.3 81–90 (part 9 of 11)

  1. Choose the correct answer from the given four options in each of the exercises from $\displaystyle 56$ to 82.

    Exercise 81

    Eight coins are tossed together. The probability of getting exactly 3\displaystyle 3 heads is
    (A)
    1256\displaystyle \frac{1}{256}
    (B)
    732\displaystyle \frac{7}{32}
    (C)
    532\displaystyle \frac{5}{32}
    (D)
    332\displaystyle \frac{3}{32}

    Matches the book, not yet reviewed

    This working reaches the answer NCERT prints. It has not yet been read through by hand.

    NCERT’s answer
    B
    (B) \(\displaystyle \frac{7}{32}\)Binomial with \(\displaystyle n=8,\ p=\tfrac12\), exactly $\displaystyle 3$ heads.\[P(X=3) = {}^8\mathrm{C}_3\left(\tfrac12\right)^3\left(\tfrac12\right)^5 \] \[= \frac{56}{256} \] \[= \frac{7}{32} \]
  2. Exercise 82

    Two dice are thrown. If it is known that the sum of numbers on the dice was less than 6\displaystyle 6, the probability of getting a sum 3\displaystyle 3, is
    (A)
    118\displaystyle \frac{1}{18}
    (B)
    518\displaystyle \frac{5}{18}
    (C)
    15\displaystyle \frac{1}{5}
    (D)
    25\displaystyle \frac{2}{5}

    Matches the book, not yet reviewed

    This working reaches the answer NCERT prints. It has not yet been read through by hand.

    NCERT’s answer
    C
    (C) \(\displaystyle \frac{1}{5}\)Reduced sample space: sums $\displaystyle 2$, $\displaystyle 3$, $\displaystyle 4$, 5.\[n(\text{sum}<6) = 1+2+3+4 = 10 \] \[n(\text{sum}=3) = 2 \quad \{(1,2),(2,1)\} \] \[P(\text{sum}=3 \mid \text{sum}<6) = \frac{2}{10} = \frac{1}{5} \]
  3. Exercise 83

    Which one is not a requirement of a binomial distribution?
    (A)
    There are 2\displaystyle 2 outcomes for each trial
    (B)
    There is a fixed number of trials
    (C)
    The outcomes must be dependent on each other
    (D)
    The probability of success must be the same for all the trials

    Matches the book, not yet reviewed

    This working reaches the answer NCERT prints. It has not yet been read through by hand.

    NCERT’s answer
    C
    (C) The outcomes must be dependent on each otherBinomial trials are independent, with two outcomes, a fixed number of trials and constant \(\displaystyle p\). Dependence is not a requirement.
  4. Exercise 84

    Two cards are drawn from a well shuffled deck of 52\displaystyle 52 playing cards with replacement. The probability, that both cards are queens, is
    (A)
    113×113\displaystyle \frac{1}{13} \times \frac{1}{13}
    (B)
    113+113\displaystyle \frac{1}{13}+\frac{1}{13}
    (C)
    113×117\displaystyle \frac{1}{13} \times \frac{1}{17}
    (D)
    113×451\displaystyle \frac{1}{13} \times \frac{4}{51}

    Matches the book, not yet reviewed

    This working reaches the answer NCERT prints. It has not yet been read through by hand.

    NCERT’s answer
    A
    (A) \(\displaystyle \frac{1}{13}\times\frac{1}{13}\)With replacement, the draws are independent.\[P(\text{queen}) = \frac{4}{52} = \frac{1}{13} \] \[P(\text{both queens}) = \frac{1}{13}\times\frac{1}{13} \]
  5. Exercise 85

    The probability of guessing correctly at least 8\displaystyle 8 out of 10\displaystyle 10 answers on a true-false type examination is
    (A)
    764\displaystyle \frac{7}{64}
    (B)
    7128\displaystyle \frac{7}{128}
    (C)
    451024\displaystyle \frac{45}{1024}
    (D)
    741\displaystyle \frac{7}{41}

    Matches the book, not yet reviewed

    This working reaches the answer NCERT prints. It has not yet been read through by hand.

    NCERT’s answer
    B
    (B) \(\displaystyle \frac{7}{128}\)Binomial with \(\displaystyle n=10,\ p=\tfrac12\); each term has \(\displaystyle \left(\tfrac12\right)^{10}\).\[P(X\ge 8) = P(8)+P(9)+P(10) \] \[= \left({}^{10}\mathrm{C}_8+{}^{10}\mathrm{C}_9+{}^{10}\mathrm{C}_{10}\right)\frac{1}{1024} \] \[= \frac{45+10+1}{1024} = \frac{56}{1024} = \frac{7}{128} \]
  6. Exercise 86

    The probability that a person is not a swimmer is 0.3. The probability that out of 5\displaystyle 5 persons 4\displaystyle 4 are swimmers is
    (A)
    5C4(0.7)4(0.3)\displaystyle { }^5 \mathrm{C}_4(0.7)^4(0.3)
    (B)
    5C1(0.7)(0.3)4\displaystyle { }^5 \mathrm{C}_1(0.7)(0.3)^4
    (C)
    5C4(0.7)(0.3)4\displaystyle { }^5 \mathrm{C}_4(0.7)(0.3)^4
    (D)
    (0.7)4(0.3)\displaystyle (0.7)^4(0.3)

    Matches the book, not yet reviewed

    This working reaches the answer NCERT prints. It has not yet been read through by hand.

    NCERT’s answer
    A
    (A) \(\displaystyle {}^5\mathrm{C}_4(0.7)^4(0.3)\)\[p = P(\text{swimmer}) = 1-0.3 = 0.7,\quad q = 0.3 \] \[P(X=4) = {}^5\mathrm{C}_4\,p^4q^{1} \] \[= {}^5\mathrm{C}_4(0.7)^4(0.3) \]
  7. Exercise 87

    The probability distribution of a discrete random variable X is given below:
    X2\displaystyle 23\displaystyle 34\displaystyle 45\displaystyle 5
    P(X)5k\displaystyle \frac{5}{k}7k\displaystyle \frac{7}{k}9k\displaystyle \frac{9}{k}11k\displaystyle \frac{11}{k}
    The value of k\displaystyle k is
    (A)
    8\displaystyle 8 (B) 16\displaystyle 16
    (C)
    32\displaystyle 32
    (D)
    48\displaystyle 48

    Matches the book, not yet reviewed

    This working reaches the answer NCERT prints. It has not yet been read through by hand.

    NCERT’s answer
    C
    (C) \(\displaystyle 32\)Probabilities of a distribution add to 1.\[\frac{5}{k}+\frac{7}{k}+\frac{9}{k}+\frac{11}{k} = 1 \] \[\frac{32}{k} = 1 \] \[k = 32 \]
  8. Exercise 88

    For the following probability distribution:
    X−4\displaystyle -4−3\displaystyle -3−2\displaystyle -2−1\displaystyle -10\displaystyle 0
    P(X)0.1\displaystyle 0.10.2\displaystyle 0.20.3\displaystyle 0.30.2\displaystyle 0.20.2\displaystyle 0.2
    E(X)\displaystyle \mathrm{E}(\mathrm{X}) is equal to :
    (A)
    0\displaystyle 0 (B) −1\displaystyle -1 (C) −2\displaystyle -2 (D) −1.8\displaystyle -1.8

    Matches the book, not yet reviewed

    This working reaches the answer NCERT prints. It has not yet been read through by hand.

    NCERT’s answer
    D
    (D) \(\displaystyle -1.8\)\[E(X) = \sum x_i\,p_i \] \[= (-4)(0.1)+(-3)(0.2)+(-2)(0.3)+(-1)(0.2)+0(0.2) \] \[= -0.4-0.6-0.6-0.2+0 \] \[= -1.8 \]
  9. Exercise 89

    For the following probability distribution
    X1\displaystyle 12\displaystyle 23\displaystyle 34\displaystyle 4
    P (X)110\displaystyle \frac{1}{10}15\displaystyle \frac{1}{5}310\displaystyle \frac{3}{10}25\displaystyle \frac{2}{5}
    E(X2)\displaystyle \mathrm{E}\left(\mathrm{X}^2\right) is equal to
    (A)
    3\displaystyle 3 (B) 5\displaystyle 5 (C) 7\displaystyle 7 (D) 10\displaystyle 10

    Matches the book, not yet reviewed

    This working reaches the answer NCERT prints. It has not yet been read through by hand.

    NCERT’s answer
    D
    (D) \(\displaystyle 10\)\[E(X^2)=\sum x^2\,P(x) \] \[E(X^2)=1^2\cdot\tfrac{1}{10}+2^2\cdot\tfrac{1}{5}+3^2\cdot\tfrac{3}{10}+4^2\cdot\tfrac{2}{5} \] \[=\tfrac{1}{10}+\tfrac{8}{10}+\tfrac{27}{10}+\tfrac{64}{10} \] \[=\tfrac{100}{10}=10 \]
  10. Exercise 90

    Suppose a random variable X follows the binomial distribution with parameters n\displaystyle n and p\displaystyle p, where 0<p<1\displaystyle 0<p<1. If P(x=r)/P(x=n−r)\displaystyle \mathrm{P}(x=r) / \mathrm{P}(x=n-r) is independent of n\displaystyle n and r\displaystyle r, then p\displaystyle p equals
    (A)
    12\displaystyle \frac{1}{2}
    (B)
    13\displaystyle \frac{1}{3}
    (C)
    15\displaystyle \frac{1}{5}
    (D)
    17\displaystyle \frac{1}{7}

    Matches the book, not yet reviewed

    This working reaches the answer NCERT prints. It has not yet been read through by hand.

    NCERT’s answer
    A
    (A) \(\displaystyle \tfrac12\)With \(\displaystyle q=1-p\): \[P(x=r)=\binom{n}{r}p^{r}q^{\,n-r},\qquad P(x=n-r)=\binom{n}{n-r}p^{\,n-r}q^{r} \] \[\binom{n}{r}=\binom{n}{n-r}\ \Rightarrow\ \frac{P(x=r)}{P(x=n-r)}=\frac{p^{r}q^{\,n-r}}{p^{\,n-r}q^{r}}=\left(\frac{p}{q}\right)^{2r-n} \] The ratio is free of \(\displaystyle n\) and \(\displaystyle r\) only when its base is $\displaystyle 1$: \[\frac{p}{q}=1\ \Rightarrow\ p=1-p\ \Rightarrow\ p=\tfrac12 \]